Integration by parts

1. Integral of x^3 e^4x dx
2. Integral of 2x E^2x dx
1
Expert's answer
2019-12-11T10:31:04-0500

1."\int x^3e^{4x}dx=x^3e^{4x}/4-(3/4)\int x^2e^{4x}dx"

"=x^3e^{4x}/4-3/4(x^2e^{4x}/4-2/4\int xe^{4x})"

"=x^3e^{4x}\/4-3\/4(x^2e^{4x}\/4-2\/4(xe^{4x}\/4-e^{4x}\/16)+c"

"=x^3e^{4x}\/4-3x^2e^{4x}\/16+(3\/32)xe^{4x}-3e^{4x}\/128)+c"

2.

"\int2xe^{2x}dx=2xe^{2x}/2-\int e^{2x}dx=xe^{2x}-e^{2x}/2+c"